OpenStudy (anonymous):

What are the coordinates of the x-intercepts of the graph of y = 2x2 + 6x – 20? (–5, 0), (2, 0) (5, 0), (–2, 0) (–4, 0), (10, 0) (4, 0), (–10, 0)

OpenStudy (anonymous):

For the x-intercepts y will be zero right? All you have to do is plug in y=0 for y = 2x2 + 6x – 20 and solve for x 0= 2x2 + 6x – 20 You might need to use the quadratic formula to solve for x.

OpenStudy (anonymous):

After you find your two x : x1 and x2 you pair it with y=0 (x1,0) & (x2,0)

OpenStudy (anonymous):

You think you can solve it with the information I gave you?

OpenStudy (anonymous):

check factoring first... 2x^2 + 6x -20 = 2 ( x^2 + 3x - 10) = 2 (x+5) (x -2)

OpenStudy (anonymous):

is it the second one?

OpenStudy (anonymous):

no, it's the first one... be careful... because one factor is (x+5), x has to be -5 so that (x+5) will be = 0. same with the other factor...

OpenStudy (anonymous):